if 1/x^2-25=1/k[1/x-5-1/x+5] then find the value of k
 

Grade:8

2 Answers

Arun
4 years ago
Dear student
 
Pleas ask question properly as it is not clear where bracket is used and which term is in numerator and denominator. Please attach an image if possible.
 
I will be happy to help you
Aditya Gupta
4 years ago
hello mrinal, the ques can simply be solved as follows:
1/(x^2-25) = 1/k[1/(x-5) – 1/(x+5)] 
or 1/(x^2-25) = 1/k[{(x+5) – (x – 5)}/(x-5)(x+5)]
or 1/(x^2-25) = 1/k[10/(x^2-25)]
or k = 10
kindly approve :))